he resistance of a wire is 10 ohm. If it is streteched ten times, the resistance will read
A) 1 ohm B) 10 ohm C) 100 ohm D) 1000ohm 
pls explain!
THANKS!

 
 
 
 

Grade:

2 Answers

Arun
4 years ago
Dear student
 
Aswe know
 
R = (rho) L/A
 
Now as L’ = 10 L
 
hence new resistance will be 10 times of previous one = 10 * 10 = 100 
Niranjan Pandey of Cambridge
4 years ago
Let the resistivity of wire be p.
Let the length of wire be l.
Let the area of cross-section of wire be A.
Then the resistence of wire is given by R=p(l/A) 
 
Assuming the wire to be cylindrical, the new length of wire br 10(l).
The voume of wire remains constant.
Let the new area of cross-section be A2.
 
Now,πr2h=volume of cylinder.
Also,πr2  =area
Thus area*length=volume.
A/q A*l=A2*(10 l)
      A=10(A2)
      A2=A/10
Thus new resistence R2= p(10l)/A2
                                  R2=p(10l)/(A/10)
                                  R2= p(100l)A                     
                                                =100(pl/A)                         
Comparing it with original resistance  R=100R i.e. 100 ohm